Solve by the elimination method.
2r-7s=-23
7r+2s=52

Answer by Alan3354(69443)

Solve by the elimination method.
2r-7s=-23
7r+2s=52
------------
Eliminate s.
Multiply the 1st eqn by 2 and the 2nd eqn by 7 to make the coefficients 14.
------------
4r - 14s = -46
49r+ 14s = 364
----------------- add
53r = 318
r = 6
------
Sub for r in either eqn and find s.
7r+2s=52
42 + 2s = 52
2s = 10
s = 5
--------
Nothing to it. You need to learn how to do this.